# HiPreNets: High-Precision Neural Networks through Progressive Training

### v1

@@ 2.1 Algorithmic Formulation of HiPreNet Training
Dataset $(\textbf{x},\textbf{f})=\{(x_{k},f(x_{k})))\}^{N}_{k=1}$, trained neural network $f_{0}^{NN}$, tolerance $tol$
@@
$r_{0}\leftarrow\textbf{f}-f_{0}^{NN}(\textbf{x})$
$e_{0}\leftarrow\max{|r_{0}|}$triangleright estimate of true $e_{0}$
@@
     $e_{i}\leftarrow e_{i-1}\cdot\max{|r_{i}(\textbf{x})|}$triangleright estimate of true $e_{i}$
@@ 2.3 Comparison with Kolmogorov–Arnold Networks (KANs)
Function Initial RMSE Final RMSE KAN RMSE Parameters per Training Total Model Size KAN Model Size
@@ 3 Network Width Adaptation in Residual Stages
Function Network Structure Final RMSE Final $L^{\infty}$ Norm Error Parameters per Training Model Size
@@ 5 Weighted Sampling Based on Residual Errors
Dataset $(\textbf{x},\textbf{f})=\{(x_{k},f(x_{k})))\}^{N}_{k=1}$, trained neural network $f_{0}^{NN}$, tolerance $tol$
@@
$e_{0}\leftarrow\max{|r_{0}|}$triangleright estimate of true $e_{0}$
@@
     $e_{i}\leftarrow e_{i-1}\cdot\max{|r_{i}(\textbf{x})|}$triangleright estimate of true $e_{i}$
@@ 6 Targeted Model Correction with Local Residual Patching
Data Initial RMSE Initial $L^{\infty}$ Norm Error Patching RMSE Patching $L^{\infty}$ Norm Error
@@ 7 Improving Boundary Accuracy through Domain Expansion
We evaluate this approach on Function I.13.12, where 1,000,000 validation samples are generated with all input variables in $[1.1,4.9]$. Two residual ensembles are trained: one on 1,000,000 samples from the domain $[1.1,4.9]$ (the baseline), and another on the slightly expanded domain $[1,5]$. We compare each ensemble’s performance on the validation data in Table 9 and in Figure 12, where we see that the ensemble trained on an expanded domain sees a substantial reduction in RMSE and $L^{\infty}$ norm error.
Function Training Domain Final RMSE on [1.1,4.9] Final $L^{\infty}$ Norm Error on [1.1,4.9]
@@
Figure 12: Validation results on data generated in $[1.1,4.9]$ at each iteration. The normal domain network is trained on data generated in $[1.1,4.9]$, while the expanded domain network is trained on data generated in $[1,5]$.
To explore the limits of this approach, we conduct a second experiment with a more pronounced domain expansion. Now, the 1,000,000 validation data samples are all generated to have input variables in $[2,4]$. Again, two residual ensembles are trained with 1,000,000 samples: one on $[2,4]$ and the other on the same expanded domain $[1,5]$. The final RMSE and $L^{\infty}$ norm error for each ensemble is given in Table 10. We now observe that the expanded domain ensemble performs noticeably worse in RMSE than the original domain ensemble. This suggests that when the expanded region is too large relative to the target domain, the model may underfit the region of interest unless given a proportionally larger training set.
Function Training Domain Final RMSE on [2,4] Final $L^{\infty}$ Norm Error on [2,4]
@@
Figure 13: Validation results on data generated in $[2,4]$ at each iteration. The normal domain network is trained on data generated in $[2,4]$, while the expanded domain network is trained on data generated in $[1,5]$.

### v2

@@ 2.1 Algorithmic Formulation of HiPreNet Training
Dataset $(\textbf{x},\textbf{f})=\{(x_{k},f(x_{k}))\}^{N}_{k=1}$, trained neural network $f_{0}^{NN}$, tolerance $tol$
@@
$r_{0}(\textbf{x})\leftarrow\textbf{f}-f_{0}^{NN}(\textbf{x})$
$e_{0}\leftarrow\max{|r_{0}(\textbf{x})|}$triangleright estimate of true $e_{0}$
@@
  $e_{i}\leftarrow\max{|r_{i}(\textbf{x})|}$triangleright estimate of true $e_{i}$
@@ 2.3 Comparison with Kolmogorov–Arnold Networks (KANs)
Function
Initial
RMSE
Final
RMSE
KAN
RMSE
Parameters
per Training
Total
Model Size
KAN
Model Size
@@ 3 Network Width Adaptation in Residual Stages
Function
Network
Structure
Final
RMSE
Final $L^{\infty}$
Norm Error
Parameters
per Training
Model
Size
@@ 5 Weighted Sampling Based on Residual Errors
Dataset $(\textbf{x},\textbf{f})=\{(x_{k},f(x_{k}))\}^{N}_{k=1}$, trained neural network $f_{0}^{NN}$, tolerance $tol$
@@
$e_{0}\leftarrow\max{|r_{0}(\textbf{x})|}$triangleright estimate of true $e_{0}$
@@
  $e_{i}\leftarrow\max{|r_{i}(\textbf{x})|}$triangleright estimate of true $e_{i}$
@@ 6 Targeted Model Correction with Local Residual Patching
Data
Initial
RMSE
Initial $L^{\infty}$
Norm Error
Patching
RMSE
Patching $L^{\infty}$
Norm Error
@@ 7 Improving Boundary Accuracy through Domain Expansion
We evaluate this approach on Function I.13.12, where 1,000,000 validation samples are generated with all input variables in [1.1,4.9][1.1,4.9][ 1.1 , 4.9 ]. Two residual ensembles are trained: one on 1,000,000 samples from the domain [1.1,4.9][1.1,4.9][ 1.1 , 4.9 ] (the baseline), and another on the slightly expanded domain [1,5][1,5][ 1 , 5 ]. We compare each ensemble’s performance on the validation data in Table 9 and in Figure 12, where we see that the ensemble trained on an expanded domain sees a substantial reduction in RMSE and $L^{\infty}$ norm error.
Function Training Domain
Final RMSE
on [1.1,4.9]
Final $L^{\infty}$ Norm Error
on [1.1,4.9]
@@
Figure 12: Validation results on data generated in [1.1,4.9][1.1,4.9][ 1.1 , 4.9 ] at each iteration. The normal domain network is trained on data generated in [1.1,4.9][1.1,4.9][ 1.1 , 4.9 ], while the expanded domain network is trained on data generated in [1,5][1,5][ 1 , 5 ].
To explore the limits of this approach, we conduct a second experiment with a more pronounced domain expansion. Now, the 1,000,000 validation data samples are all generated to have input variables in [2,4][2,4][ 2 , 4 ]. Again, two residual ensembles are trained with 1,000,000 samples: one on [2,4][2,4][ 2 , 4 ] and the other on the same expanded domain [1,5][1,5][ 1 , 5 ]. The final RMSE and $L^{\infty}$ norm error for each ensemble is given in Table 10. We now observe that the expanded domain ensemble performs noticeably worse in RMSE than the original domain ensemble. This suggests that when the expanded region is too large relative to the target domain, the model may underfit the region of interest unless given a proportionally larger training set.
Function Training Domain
Final RMSE
on [2,4]
Final $L^{\infty}$ Norm Error
on [2,4]
@@
Figure 13: Validation results on data generated in [2,4][2,4][ 2 , 4 ] at each iteration. The normal domain network is trained on data generated in [2,4][2,4][ 2 , 4 ], while the expanded domain network is trained on data generated in [1,5][1,5][ 1 , 5 ].